# Enzyme-Substrate Complex Formation Modulates Diffusion-Driven Patterning in Metabolic Pathways

### v1

@@ Abstract.
We investigate how enzymatic binding kinetics regulate diffusion-driven instabilities in a two-step metabolic pathway. Starting from a mechanistic description in which the substrate reversibly binds to the first enzyme before catalytic conversion, we formulate two reaction–diffusion models: a simplified system with effective kinetics and a non-degenerated system that explicitly includes the transient enzyme–substrate complex. A quasi-steady-state reduction yields a lower-dimensional model that preserves the nonlinear coupling between catalytic turnover and spatial transport while enabling complete linear and weakly nonlinear analyses. We show that explicit complex formation shifts the homogeneous steady state, modifies relaxation dynamics, and substantially alters the size and location of the Turing instability region relative to the simplified model. Numerical simulations agree closely with weakly nonlinear predictions, demonstrating how reversible binding attenuates pattern amplitude and slows the emergence of spatial heterogeneity. These results establish a quantitative link between enzyme–substrate binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as that associated with liquid–liquid phase separation (LLPS). The framework provides a mechanistic route by which biochemical association, dissociation, and catalytic rates jointly regulate the robustness and structure of spatial metabolic patterns, and can be extended to broader classes of compartmentalized biochemical networks.
@@ 1. Introduction
Understanding the complex interactions in biochemical reaction networks is essential for uncovering the principles of cellular function and metabolic regulation. Among the key features of these systems are reaction-diffusion processes, in which chemical species diffuse through space and interact in nonlinear ways. These interactions can give rise to striking spatial organizations, most notably Turing patterns, which are periodic structures that emerge spontaneously from homogeneous initial conditions when reaction kinetics and diffusion rates satisfy certain instability criteria. Since the pioneering work of Turing [37], reaction-diffusion mechanisms provide a robust mathematical framework for the generation of spatial patterns in biological systems [30, 29]. ranging from chemical oscillators [9], [11] to developmental patterning in multicellular systems [10, 15], .
Metabolic pathways in living organisms provide a particularly rich setting for studying such processes. These pathways consist of sequences of enzymatic reactions that convert substrates into intermediates and final products, often with multiple branches and feedback loops [16], [24], [17]. Spatial aspects of metabolism have received growing attention in recent years, as it has become clear that enzymes and metabolites are not always homogeneously distributed in the cytoplasm. Instead, cells frequently organize enzymes into clusters or into biomolecular condensates through liquid–liquid phase separation (LLPS) [31], [25], [7]. Such mesoscale organization can accelerate metabolic fluxes, buffer fluctuations, and regulate competition at branch points [3], [28], [26], [23, 27]. Liquid‒liquid phase separation is a fundamental mechanism of sepsis [6] From a theoretical perspective, reaction–diffusion models offer a natural framework to investigate how enzyme clustering and LLPS influence the spatial dynamics of metabolism [5, 32], [4, 8, 34, 35, 22]. For instance, liquid-liquid phase separation (LLPS) on cell membranes has been described using reaction-diffusion models, which demonstrate that enzymatic feedback loops may drive out-of-equilibrium dynamics that replicate classical phase separation properties as nucleation, coarsening, and domain creation [35]. A straightforward reaction-diffusion framework called the swarm model has been put out to explain the molecular condensation processes that occur in cells. It demonstrates how these condensates emerge out of equilibrium and may be controlled by nucleation and cooperativity dynamics. It captures important aspects of membraneless organelle development, including nucleation, clustering, diffusion, and particle exchange [8].
@@
The novelty of this work lies in combining mechanistic enzyme–substrate kinetics with diffusion and cross-diffusion to study the emergence of spatial metabolic organization. We introduce a non-degenerated reaction–diffusion formulation that explicitly incorporates reversible enzyme–substrate complex formation and derive a reduced model via a quasi-steady-state approximation that preserves the essential nonlinear feedback between catalysis and spatial transport. By comparing this model with a classical simplified enzymatic pathway, we demonstrate how complex formation shifts the homogeneous steady state, alters relaxation dynamics, and significantly modifies the Turing instability region. To our knowledge, this is the first analysis quantifying how enzymatic binding kinetics reshape diffusion-driven instabilities in two-step metabolic pathways, providing a mechanistic link between enzyme-mediated feedback and mesoscale organization such as liquid–liquid phase separation.
@@
As a starting point, we consider a simplified two-step metabolic pathway [5]: $S_{0}\xrightarrow{E_{1}}S_{1}\xrightarrow{E_{2}}P,$ where $S_{0}$, $S_{1}$, and $P$ denote substrate, intermediate, and final product, respectively. The substrate $S_{0}$ is converted to the intermediate $S_{1}$ by enzyme $E_{1}$, which in turn is converted to $P$ by enzyme $E_{2}$. Following the phenomenological formulation of Castellana et al. [5], we model the spatio-temporal evolution of the substrate and intermediate as
@@
Such curvature-dependent transport is consistent with mesoscale biophysical processes. Curvature-sensing proteins in membranes preferentially bind to negatively curved regions [1], and curvature gradients can drive molecular fluxes in cells and tissues [36]. Analogous behavior occurs in LLPS condensates and metabolon-like assemblies, where spatial variations in viscosity, binding affinity, and chemical potential bias metabolite transport [5, 38]. The term $d\,C_{0}\nabla^{2}C_{1}$ therefore provides a minimal and biophysically motivated representation of curvature-driven substrate recruitment.
LLPS proceeds through nucleation, maturation, and functional biochemical activity [19, 2]. Model 1 corresponds to the latter two stages: we assume that a condensate has already formed and maintains approximately constant internal enzyme concentrations $n_{1}$ and $n_{2}$, while the metabolites $C_{0}$ and $C_{1}$ remain mobile and diffuse into, out of, and within the condensed phase. The curvature-driven term $d\,C_{0}\nabla^{2}C_{1}$ captures substrate recruitment into regions of high intermediate accumulation, a hallmark of the mature, functional stage of LLPS-mediated metabolic organization.
@@
Model 1 captures basic spatial coupling but treats catalysis as an effective one-step process and does not resolve the transient enzyme–substrate complex. In cellular environments, however, enzymes bind substrates reversibly before turnover [20], and the mobility of such complexes—especially inside co-localized assemblies such as liquid–liquid phase-separated (LLPS) condensates—may differ substantially from that of free metabolites.
@@
Remark. Equation (1.2) is not intended as a strict microscopic mass-action description. Instead, $C_{E_{1}S_{0}}$ acts as an effective complex-density field summarizing local co-localization of enzyme and substrate—for example within LLPS condensates or metabolon clusters. The terms $k_{a_{1}}n_{1}C_{0}$ and $k_{d_{1}}n_{1}C_{E_{1}S_{0}}$ therefore represent coarse-grained recruitment and release processes characteristic of mesoscale organization [5, 38].
@@
While the explicit inclusion of enzyme–substrate complexes enriches the model biologically, it introduces mathematical complications. In particular, the extended three-variable system is degenerate with respect to weakly nonlinear (WNL) analysis, preventing the direct derivation of amplitude equations near the instability threshold. This difficulty arises from the presence of fast binding-unbinding dynamics that render the model stiff and obstruct the perturbative expansion required in WNL theory. To overcome this challenge, we apply a quasi-steady-state reduction by assuming that the enzyme-substrate complex equilibrates rapidly relative to the other variables. This reduction yields a non-degenerate two-variable model (in section 3.2) in which the influence of the complex is encoded in effective reaction terms. Importantly, the reduced model preserves the essential nonlinear feedbacks and LLPS-related cross-diffusion couplings of the modified model, while restoring analytical tractability.
@@
Although the numerical simulations presented in Section 6 are performed in one spatial dimension, the reaction–diffusion mechanism is not restricted to 1D. The same chemical interactions and diffusion laws apply in two and three dimensions, and the Turing conditions are dimension-independent: what changes is only the geometry of the domain and the admissible wave numbers. In higher dimensions, additional pattern morphologies (spots, stripes, labyrinths, hexagons)([33], [13, 14] may occur, but the biochemical interpretation—enzymatic feedback coupled with substrate/intermediate mobility—remains unchanged. Thus, the 1D setting captures the essential instability mechanism while allowing clearer analytical and numerical comparison between the simplified and non-degenerated models.
The remainder of this paper is organized as follows. In Section 2, we analyze the local dynamics of the primary model and derive the conditions for Turing instability driven by diffusion and cross-diffusion effects. In Section 3, we examine the equilibrium structure of the extended three-variable system, showing that it is degenerate. To address this, we employ a quasi–steady-state approximation [20, 33] to obtain a reduced non-degenerate model, for which we investigate both local stability and the onset of Turing instability. Section 4 explores how variations in reaction and diffusion parameters affect the extent of Turing regions in Models 1 and 3. In Section 5, we apply weakly nonlinear (WNL) analysis to the primary and reduced models to derive amplitude equations describing the nonlinear evolution of spatial patterns. These analytical predictions are then compared with numerical simulations in Section 6, confirming the validity of the theoretical results. Finally, Section 7 presents concluding remarks and discusses the broader implications of our findings for understanding spatial regulation in enzyme-mediated and phase-separated biochemical systems.
@@ 2.1. Turing Analysis
$\mathcal{D}(E^{*})=\begin{pmatrix}D_{0}&&-dC_{0}^{E}\\ 0&&D_{1}\end{pmatrix}$ (2.3)
@@
3. Modified Model and Quasi–Steady–State Reduction
@@ 3.1. Equilibria and Stability Analysis of the Modified Model
To obtain a more realistic representation of the enzyme–substrate interaction within the reaction–diffusion framework, it is necessary to explicitly account for the formation of enzyme–substrate (intermediate) complexes. In enzymatic reactions, the catalytic conversion of the substrate into the product occurs only after the formation of such a complex, which is typically assumed to be irreversible. Therefore, the reaction scheme has been extended by including the additional binding and catalytic steps, leading to the formulation of the modified model (1.2).
We first analyze the equilibria of system (1.2) in the spatially homogeneous case. The system admits two distinct sets of steady states:
  • 1.
@@
  • 2.
    Supposing $\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}C_{E_{1}S_{0}}^{*}>0$, the family of equilibria
    $E_{2}^{*}=\Big(\dfrac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}C_{E_{1}S_{0}}^{*}}{\alpha_{0}},\,C_{E_{1}S_{0}}^{*},\,\dfrac{k_{cat}C_{E_{1}S_{0}}^{*}}{k_{2}n_{2}+\beta_{2}}\Big),$
    which appears in the degenerate case $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ and corresponds to a continuous line of equilibria.
Indeed, the degeneracy arises because the binding/unbinding dynamics are fast relative to metabolite dynamics, giving one zero eigenvalue in the Jacobian and producing a line of equilibria. This prevents standard WNL expansion, motivating the QSSA.
@@
Let $E_{1}^{*}$ denote the equilibrium point of system (1.2). Then:
@@
    $E_{1}^{*}$ is locally asymptotically stable if $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})<0$.
@@
Linearizing system (1.2) around the equilibrium $E_{1}^{*}$ yields the Jacobian matrix
$J(E_{1}^{*})=\begin{bmatrix}-\alpha_{0}&-k_{d_{1}}n_{1}&0\\[5.69054pt] 0&k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})&0\\[5.69054pt] 0&k_{cat}&-(k_{2}n_{2}+\beta_{2})\end{bmatrix}$. (3.1)
The eigenvalues of $J(E_{1}^{*})$ are $\lambda_{1}=-\alpha_{0}$, $\lambda_{2}=-(k_{2}n_{2}+\beta_{2})$, and $\lambda_{3}=k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})$. Since $\alpha_{0}>0$ and $(k_{2}n_{2}+\beta_{2})>0$, the first two eigenvalues are negative. The sign of $\lambda_{3}$ therefore determines the local stability of $E_{1}^{*}$. If $\lambda_{3}<0$, all eigenvalues are negative and the equilibrium is locally asymptotically stable. Conversely, if $\lambda_{3}>0$, one eigenvalue is positive and $E_{1}^{*}$ becomes unstable. This completes the proof. ∎
The above theorem shows that the stability of the equilibrium $E_{1}^{*}$ in the modified (non-degenerated) enzyme model depends solely on the relative magnitudes of the association, dissociation, and catalytic rates. When the binding rate $k_{a_{1}}n_{1}$ exceeds the sum of the dissociation and catalytic rates $(k_{d_{1}}+k_{cat})$, the feedback becomes self-amplifying, leading to an unstable steady state. Conversely, when the effective dissociation and catalytic losses dominate, the equilibrium remains locally asymptotically stable.
Remark 3.2.
At the critical value $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$, the Jacobian matrix possesses two negative eigenvalues and one zero eigenvalue, and the system exhibits a one-dimensional manifold of equilibria. Each point on this line is non-hyperbolic and neutrally stable along the direction of the equilibrium manifold.
Theorem 3.3.
Degenerate case: line of equilibria and neutral direction] Assume $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ in the modified model (1.2). Then:
1. The set $\mathcal{M}\;=\;\Big\{\,E_{2}^{*}(\xi)\;=\;\Big(\frac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}\xi}{\alpha_{0}}\;,\;\xi\;,\;\frac{k_{cat}}{k_{2}n_{2}+\beta_{2}}\,\xi\Big)\;:\;\xi\in\mathcal{I}\Big\}$ (with $\mathcal{I}$ chosen to enforce nonnegativity) is a one–dimensional manifold of equilibria.
2. For every $E\in\mathcal{M}$, the Jacobian $J(E)$ has eigenvalues $\lambda_{1}=-\alpha_{0}<0,\;\lambda_{2}=-(k_{2}n_{2}+\beta_{2})<0,\;\lambda_{3}=0.$ In particular, each $E\in\mathcal{M}$ is non-hyperbolic, with a neutral direction tangent to $\mathcal{M}$ and exponential stability in the two transverse directions.
Proof.
Setting the time derivatives in system (1.2) to zero and imposing the degeneracy condition $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$ removes one scalar constraint among the steady–state equations. Solving the resulting algebraic system yields the one–parameter family of equilibria $E_{2}^{*}(\xi)$ described in (1); hence every point on $\mathcal{M}$ is an equilibrium.
Evaluating the Jacobian matrix (3.1) at the degenerate condition $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$ and at $E_{2}^{*}(\xi)$ gives three eigenvalues: $\lambda_{1}=-\alpha_{0}<0$, $\lambda_{2}=-(k_{2}n_{2}+\beta)<0$, and $\lambda_{3}=0$. A tangent vector to $\mathcal{M}$ at $E_{2}^{*}(\xi)$ is $v=\frac{d}{d\xi}E_{2}^{*}(\xi)=\Big(-\tfrac{k_{d_{1}}n_{1}}{\alpha_{0}},\,1,\,\tfrac{k_{cat}}{k_{2}n_{2}+\beta}\Big)^{\!\top}.$ Direct substitution shows that $J(E_{2}^{*}(\xi))\,v=0$, so $v$ spans the nullspace of the Jacobian, confirming that the zero eigenvalue corresponds to the direction tangent to the equilibrium manifold $\mathcal{M}$.
According to the Center Manifold Theorem ([21]), a zero eigenvalue of the Jacobian gives rise to a one-dimensional local center manifold $W^{c}_{\mathrm{loc}}(E)$ tangent to the associated eigenvector. In the present case, this manifold coincides locally with the equilibrium line $\mathcal{M}$. The remaining two eigenvalues are strictly negative, and by the Stable Manifold Theorem, they generate a two–dimensional local stable manifold $W^{s}_{\mathrm{loc}}(E)$ on which perturbations decay exponentially to $\mathcal{M}$. Consequently, each $E\in\mathcal{M}$ is a non–hyperbolic equilibrium, characterized by exponential contraction in the directions transverse to $\mathcal{M}$ and neutral (center) stability along the manifold itself. ∎

### v2

@@ Abstract.
We investigate how enzymatic binding kinetics regulate diffusion-driven instabilities in a two-step metabolic pathway. Starting from a mechanistic description in which the substrate reversibly binds to the first enzyme before catalytic conversion, we formulate two reaction–diffusion models: a simplified system with effective kinetics and an extended model that explicitly includes the enzyme–substrate complex. The latter exhibits a structural degeneracy at critical parameter values due to a continuous family of homogeneous equilibria. To enable direct comparison and analytical progress, we introduce a reduced non-degenerate formulation via a quasi-equilibrium closure that encodes the influence of complex formation into effective reaction terms while preserving the nonlinear coupling between catalytic turnover and spatial transport.
We show that explicit enzyme–substrate binding shifts the homogeneous steady state, modifies relaxation dynamics, and substantially alters the size and location of the Turing instability region relative to the simplified model. Numerical simulations are in close agreement with weakly nonlinear predictions, illustrating how reversible binding reshapes pattern selection and slows the development of spatial heterogeneity.
These results establish a quantitative link between enzyme–substrate binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization, including structures associated with liquid–liquid phase separation (LLPS). The proposed framework provides a mechanistic route by which association, dissociation, and catalytic rates jointly regulate the robustness and structure of spatial metabolic patterns, and can be extended to broader classes of compartmentalized biochemical networks.
@@ 1. Introduction
Understanding the complex interactions in biochemical reaction networks is essential for uncovering the principles of cellular function and metabolic regulation. Among the key features of these systems are reaction-diffusion processes, in which chemical species diffuse through space and interact in nonlinear ways. These interactions can give rise to striking spatial organizations, most notably Turing patterns, which are periodic structures that emerge spontaneously from homogeneous initial conditions when reaction kinetics and diffusion rates satisfy certain instability criteria. Since the pioneering work of Turing (Turing (1952)), reaction-diffusion mechanisms provide a robust mathematical framework for the generation of spatial patterns in biological systems (Maini et al. (2012); Maini and Woolley (2019)). ranging from chemical oscillators (Epstein et al. (2012), Gambino et al. (2015)) to developmental patterning in multicellular systems (Furusawa and Kaneko (1998); Giunta et al. (2021)).
Metabolic pathways in living organisms provide a particularly rich setting for studying such processes. These pathways consist of sequences of enzymatic reactions that convert substrates into intermediates and final products, often with multiple branches and feedback loops (Hafner and Hatzimanikatis (2021), Lee et al. (2025), Hill et al. (2015)). Spatial aspects of metabolism have received growing attention in recent years, as it has become clear that enzymes and metabolites are not always homogeneously distributed in the cytoplasm. Instead, cells frequently organize enzymes into clusters or into biomolecular condensates through liquid–liquid phase separation (LLPS) (Milicevic et al. (2022), Li et al. (2025), Chen et al. (2022)). Such mesoscale organization can accelerate metabolic fluxes, buffer fluctuations, and regulate competition at branch points (Bevilacqua et al. (2024), Lu et al. (2021), Lim and Clark (2024), Krainer et al. (2021); Liu et al. (2023)). Liquid‒liquid phase separation is a fundamental mechanism of sepsis (Chen et al. (2025)) From a theoretical perspective, reaction-diffusion models offer a natural framework to investigate how enzyme clustering and LLPS influence the spatial dynamics of metabolism (Castellana et al. (2014); Mukherjee et al. (2024), Buchner et al. (2013); Cieza Huaman (2022); Nakasone and Terazima (2021); Rigano (2022); Kim et al. (2024)). For instance, liquid-liquid phase separation (LLPS) on cell membranes has been described using reaction-diffusion models, which demonstrate that enzymatic feedback loops may drive out-of-equilibrium dynamics that replicate classical phase separation properties as nucleation, coarsening, and domain creation Rigano (2022). A straightforward reaction-diffusion framework called the swarm model has been put out to explain the molecular condensation processes that occur in cells. It demonstrates how these condensates emerge out of equilibrium and may be controlled by nucleation and cooperativity dynamics. It captures important aspects of membraneless organelle development, including nucleation, clustering, diffusion, and particle exchange (Cieza Huaman (2022)).
@@
The novelty of this work lies in combining mechanistic enzyme–substrate kinetics with diffusion and cross-diffusion to study the emergence of spatial metabolic organization. We introduce a non-degenerated reaction–diffusion formulation that explicitly incorporates reversible enzyme–substrate complex formation and derive a reduced model via a quasi-equilibrium approximation that preserves the essential nonlinear feedback between catalysis and spatial transport. By comparing this model with a classical simplified enzymatic pathway, we demonstrate how complex formation shifts the homogeneous steady state, alters relaxation dynamics, and significantly modifies the Turing instability region. To our knowledge, this is the first analysis quantifying how enzymatic binding kinetics reshape diffusion-driven instabilities in two-step metabolic pathways, providing a mechanistic link between enzyme-mediated feedback and mesoscale organization such as liquid–liquid phase separation.
@@
As a starting point, we consider a simplified two-step metabolic pathway (Castellana et al. (2014)): $S_{0}\xrightarrow{E_{1}}S_{1}\xrightarrow{E_{2}}P,$ where $S_{0}$, $S_{1}$, and $P$ denote substrate, intermediate, and final product, respectively. The substrate $S_{0}$ is converted to the intermediate $S_{1}$ by enzyme $E_{1}$, which in turn is converted to $P$ by enzyme $E_{2}$. Following the phenomenological formulation of (Castellana et al. (2014)), we model the spatio-temporal evolution of the substrate and intermediate as
@@
Such curvature-dependent transport is consistent with mesoscale biophysical processes. Curvature-sensing proteins in membranes preferentially bind to negatively curved regions (Arnold et al. (2025)), and curvature gradients can drive molecular fluxes in cells and tissues (Schamberger et al. (2023)). Analogous behavior occurs in LLPS condensates and metabolon-like assemblies, where spatial variations in viscosity, binding affinity, and chemical potential bias metabolite transport (Castellana et al. (2014); Zwicker et al. (2017)). The term $d\,C_{0}\nabla^{2}C_{1}$ therefore provides a minimal and biophysically motivated representation of curvature-driven substrate recruitment.
LLPS proceeds through nucleation, maturation, and functional biochemical activity (Hyman et al. (2014); Banani et al. (2017)). Model 1 corresponds to the latter two stages: we assume that a condensate has already formed and maintains approximately constant internal enzyme concentrations $n_{1}$ and $n_{2}$, while the metabolites $C_{0}$ and $C_{1}$ remain mobile and diffuse into, out of, and within the condensed phase. The curvature-driven term $d\,C_{0}\nabla^{2}C_{1}$ captures substrate recruitment into regions of high intermediate accumulation, a hallmark of the mature, functional stage of LLPS-mediated metabolic organization.
@@
Model 1 captures basic spatial coupling but treats catalysis as an effective one-step process and does not resolve the transient enzyme–substrate complex. In cellular environments, however, enzymes bind substrates reversibly before turnover (Keener and Sneyd (2009)), and the mobility of such complexes—especially inside co-localized assemblies such as liquid–liquid phase-separated (LLPS) condensates—may differ substantially from that of free metabolites.
@@
Remark. Equation (1.2) is not intended as a strict microscopic mass-action description. Instead, $C_{E_{1}S_{0}}$ acts as an effective complex-density field summarizing local co-localization of enzyme and substrate—for example within LLPS condensates or metabolon clusters. The terms $k_{a_{1}}n_{1}C_{0}$ and $k_{d_{1}}n_{1}C_{E_{1}S_{0}}$ therefore represent coarse-grained recruitment and release processes characteristic of mesoscale organization (Castellana et al. (2014); Zwicker et al. (2017)).
@@
While explicitly including the enzyme–substrate complex enriches the model biologically, the homogeneous dynamics of Model 2 becomes degenerate at the critical relation $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0$, where the system admits a one-dimensional manifold of equilibria and the Jacobian has a zero eigenvalue. This non-hyperbolicity obstructs the direct application of standard weakly nonlinear (WNL) theory in its simplest form, which typically assumes an isolated equilibrium at the bifurcation point.
To enable direct comparison with the primary two-variable pathway and to restore analytical tractability, we introduce a reduced two-variable non-degenerated model (Section 3.2) by eliminating $C_{E_{1}S_{0}}$ through a quasi-equilibrium closure that neglects its diffusive contribution at leading order and expresses the complex concentration in terms of the substrate. The resulting model preserves the mechanistic influence of complex formation through modified effective reaction coefficients, while retaining the same cross-diffusion structure that encodes spatial recruitment in LLPS-like environments.
@@
Although the numerical simulations presented in Section 6 are performed in one spatial dimension, the reaction–diffusion mechanism is not restricted to 1D. The same chemical interactions and diffusion laws apply in two and three dimensions, and the Turing conditions are dimension-independent: what changes is only the geometry of the domain and the admissible wave numbers. In higher dimensions, additional pattern morphologies (spots, stripes, labyrinths, hexagons)(Murray (2007), Gambino et al. (2013, 2019)) may occur, but the biochemical interpretation—enzymatic feedback coupled with substrate/intermediate mobility—remains unchanged. Thus, the 1D setting captures the essential instability mechanism while allowing clearer analytical and numerical comparison between the simplified and non-degenerated models.
The remainder of this paper is organized as follows. In Section 2, we analyze the local dynamics of the primary model and derive the conditions for Turing instability driven by diffusion and cross-diffusion effects. In Section 3, we examine the equilibrium structure of the extended three-variable system, showing that it is degenerate. To address this, we employ a quasi-equilibrium approximation (Keener and Sneyd (2009); Murray (2007)) to obtain a reduced non-degenerate model, for which we investigate both local stability and the onset of Turing instability. Section 4 explores how variations in reaction and diffusion parameters affect the extent of Turing regions in Models 1 and 3. In Section 5, we apply weakly nonlinear (WNL) analysis to the primary and reduced models to derive amplitude equations describing the nonlinear evolution of spatial patterns. These analytical predictions are then compared with numerical simulations in Section 6, confirming the validity of the theoretical results. Finally, Section 7 presents concluding remarks and discusses the broader implications of our findings for understanding spatial regulation in enzyme-mediated and phase-separated biochemical systems.
@@ 2.1. Turing Analysis
$\mathcal{D}(E^{*})=\begin{pmatrix}D_{0}&&-dC_{0}^{E}\\ 0&&D_{1}\end{pmatrix}$, (2.3)
@@
3. Modified Model and Quasi-Equilibrium Approximation
@@ 3.1. Equilibria and Stability Analysis of the Modified Model
To obtain a more mechanistic representation of the enzymatic conversion process, the reaction scheme is extended to include the reversible formation of the enzyme–substrate complex and its catalytic conversion into the intermediate. This leads to the three–species reaction–diffusion system (1.2), which we first analyse in the spatially homogeneous setting.
Equilibria. Setting time derivatives to zero in the homogeneous system yields two types of steady states:
  • (1)
@@
  • (2)
    When the condition $k_{a_{1}}n_{1}-(\,k_{d_{1}}+k_{cat}\,)=0$ holds, the steady–state equations become linearly dependent and admit a one–parameter family of equilibria $E_{2}^{*}(\xi)=\left(\frac{\alpha_{0}C_{0}^{*}-k_{d_{1}}n_{1}\,\xi}{\alpha_{0}},\;\xi,\;\frac{k_{cat}}{k_{2}n_{2}+\beta_{2}}\,\xi\right),\qquad\xi\in\mathcal{I},$ where $\mathcal{I}$ is chosen to ensure nonnegativity of the components. This degeneracy corresponds to a continuous curve of equilibria.
Stability of $E_{1}^{*}$. Linearization of system (1.2) around $E_{1}^{*}$ gives
$J(E_{1}^{*})=\begin{bmatrix}-\alpha_{0}&-k_{d_{1}}n_{1}&0\\[2.84526pt] 0&k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})&0\\[2.84526pt] 0&k_{cat}&-(k_{2}n_{2}+\beta_{2})\end{bmatrix}$, (3.1)
whose eigenvalues are $\lambda_{1}=-\alpha_{0},\qquad\lambda_{2}=-(k_{2}n_{2}+\beta_{2}),\qquad\lambda_{3}=k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat}).$ Because $\alpha_{0}>0$ and $(k_{2}n_{2}+\beta_{2})>0$, the first two eigenvalues are strictly negative, and the sign of $\lambda_{3}$ determines the local stability of $E_{1}^{*}$.
@@
Let $E_{1}^{*}$ denote the equilibrium of system (1.2). Then:
@@
    If $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})<0$, then $E_{1}^{*}$ is locally asymptotically stable.
@@
Immediate from the eigenvalues of $J(E_{1}^{*})$, as described above. ∎
Degenerate case: line of equilibria. At the critical value $k_{a_{1}}n_{1}-(k_{d_{1}}+k_{cat})=0,$ the Jacobian at any equilibrium point $E_{2}^{*}(\xi)$ possesses two strictly negative eigenvalues and one zero eigenvalue. Thus every point on the equilibrium curve $\mathcal{M}=\{E_{2}^{*}(\xi)\}$ is non–hyperbolic: perturbations orthogonal to $\mathcal{M}$ decay exponentially, while perturbations tangent to $\mathcal{M}$ neither grow nor decay. This neutral direction reflects the structural degeneracy of the steady–state equations and the presence of a continuous family of equilibria.
Geometrically, the set $\mathcal{M}$ forms a one–dimensional invariant manifold of steady states, with exponential attraction in the two transverse directions. Perturbations therefore relax rapidly onto the neighbourhood of $\mathcal{M}$, after which the dynamics remains close to this manifold.